Circular No. 8925

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 E2 (LINEAR) The following ephemeris was computed from the orbital elements given

on IAUC 8924 (H_{10} = 18.0):

2008 TT α_{2000} δ_{2000} Δ r ϵ β Mag.
Mar. 5 16^{h}02.^{m}06 +24^{o}16.^{′}4 0.777 1.423 106.^{o}5 42.^{o}0 19.0
15 16 31.89 +23 46.6 0.669 1.347 106.5 45.1 18.4
25 17 04.70 +22 28.7 0.568 1.276 105.9 48.7 17.8
Apr. 4 17 42.05 +19 50.0 0.476 1.211 104.6 53.0 17.2
14 18 26.18 +14 57.7 0.394 1.155 102.7 57.9 16.6
24 19 20.08 + 6 42.6 0.329 1.110  99.9 63.2 16.0
May 4 20 25.70 − 5 24.6 0.290 1.079  96.2 68.3 15.6
14 21 40.17 −19 02.5 0.288 1.063  92.4 71.9 15.6
24 22 54.20 −29 49.8 0.321 1.062  90.0 72.4 15.8
June 3 23 57.84 −36 22.1 0.375 1.079  89.5 70.1 16.2
13 0 47.22 −40 01.3 0.439 1.110  90.6 66.2 16.7
23 1 23.54 −42 16.1 0.503 1.155  92.7 61.5 17.1
2001 QQ_{322} AND 2005 PR_{21}

K. S. Noll, Space Telescope Science Institute (STScI); W. M. Grundy, Lowell Observatory; S. D. Kern, STScI; H. F. Levison, Southwest Research Institute; and D. C. Stephens, Brigham Young University, report that the transneptunian objects 2001 QQ_{322} (cf. MPEC 2001-V11) and 2005 PR_{21} (cf. MPECs 2005-T100, 2006-T35) both have similar-sized companions (based on brightness). The observations of 2001 QQ_{322} were made during 2007 June 15.0197–15.0535 UT with the Planetary Camera of the Wide Field Planetary Camera 2 on the Hubble Space Telescope (HST), using the F606W filter (wide V) with one 260-s exposure at four dithered positions on the detector. The two components of 2001 QQ_{322} were separated by an angular distance of 0′′.1272 ± 0′′.0015, with the secondary being fainter by 0.2 magnitude; the secondary was located at −0′′.105 ± 0′′.003 in α and −0′′.072 ± 0′′.002 in δ relative to the primary. The observations of 2005 PR_{21} were made as above during 2007 May 10.5156–10.6019 (but only with exposures at two dithered positions on the detector). The two components of 2005 PR_{21} were separated by an angular distance of 0′′.123 ± 0′′.009, with the secondary being fainter by 1.1 magnitude; the secondary was located at −0′′.122 ± 0′′.009 in α and −0′′.015 ± 0′′.002 in δ relative to the primary.

2008 March 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green